Find the periodic payment R required to accumulate a sum of S dollars over t years with interest earned at the rate of r%/year compounded m times a year. (Round your answer to the nearest cent.) S = 40,000, r = 5, t = 6, m = 2
Answer #1

SOLUTION :


Following formula will be applicable for periodic payment $R to accumulate $S in t years at interest rate of r% per year compounded m times a year ;


S = R((1 + r / (100m))^(mt) - 1) / (r / (100m))   


For S = $40000 ; r = 5% ; t = 6 years ; m = 2 , we have :


=> 40000 = R ( (1 + 5 / 200)^(12) - 1 ) / (5 / 200)

=> 40000 = R * 13.795553

=> R = 40000/13.795553

=> R = 2899.49 ($) (ANSWER)

answered by: Tulsiram Garg